### v1

# SNIC: Synthesized Noisy Images using Calibration

SNIC: Synthesized Noisy Images using Calibration
Nik Bhatt
Columbia University
nb3097@columbia.edu
1 Abstract
Advanced denoising algorithms require large, high-quality datasets. Physics-based, statistical noise models can create such datasets by realistically simulating noise in digital images. However, there is little information on the correct way to calibrate and tune these heteroscedastic models, and a lack of published datasets using them. In this paper, we explore the process of building high-quality heteroscedastic noise models. Our methods produce realistic synthesized noisy images, in both RAW and TIFF formats. Our synthesized noisy images achieve comparable LPIPS results to real noisy images, and greatly outperform those created with manufacturer-provided DNG noise models both in LPIPS and when tested with a state-of-the-art (SOTA) denoising model. Using our approach, we created the Synthesized Noisy Images using Calibration dataset (SNIC) containing over 6000 noisy images, comprising 30 scenes from four sensors, including two smartphone sensors, a point-and-shoot, and a DSLR. SNIC is the first synthesized noisy image dataset provided in both RAW and TIFF format.
2 Introduction
All digital images contain noise as a by-product of the capture process. Noise arises from the quantum effects of photon arrival, and from various electronic components in the sensor and camera circuitry, including sensor readout, amplifiers, and analog-to-digital conversion. Images captured in low-light environments, and especially with the smaller sensors found on smartphones, typically include considerable amounts of noise, negatively affecting the appearance and aesthetics of the images. Thus, noise reduction, or denoising, is an important and difficult task that has been the focus of active research for years. Denoising is an ill-posed inverse problem with no one correct solution, but rather an infinite set of possible solutions. Denoising is best performed on RAW images (rather than camera-generated JPEGs) because RAW images are scene-referred, undergo less processing by the image signal processor (ISP), contain linear data, and do not include gamma correction.
Currently most research in denoising focuses on machine learning models. Thus, researchers developing new denoising algorithms rely on large image datasets to train and test their models. There are three main approaches to acquiring an adequate dataset.
  • First, one can add noise to images from another dataset such as Flickr2K. Because of the wide range of cameras in these datasets, the noise is typically introduced using Additive White Gaussian Noise (AWGN), a simple homoscedastic model that does not account for the way noise actually appears in digital images. Consequently, AWGN-based synthetic images lack realism, failing to capture the true noise characteristics of digital sensors.
  • Second, one can use an existing real noisy dataset that consists of low ISO and high ISO images that have been extremely carefully aligned. Making a real noise dataset is exceptionally difficult, so there are only a few available. Most were created in 2017 or 2018, and thus use older cameras such as the iPhone 7, released over nine years ago. Real noise datasets are also often limited to specially lit indoor scenes.
  • Third, one can add noise to one’s own images using a heteroscedastic Poisson-Gaussian noise model, which is superior to a homoscedastic one like AWGN. A heteroscedastic model can be easily retrieved from DNG images, though we found those models are often inaccurate and incomplete. We have chosen a variant of the third approach: create a custom physics-based heteroscedastic noise model by calibrating cameras and then tuning the model with validation images. This requires more effort than simply using DNG noise profiles, but our models provide perfectly aligned, high-quality noisy RAW images. Further, this approach offers full control over the choice of cameras and scenes used.
The choice and quality of the dataset directly affects the performance of a denoising algorithm on real images. For physically modeled datasets, the key is minimizing the discrepancy between the synthetic noise and actual noise [1]. In this paper, we focus on physics-based noise modeling, using careful calibration to determine photon shot noise, combined with complete capture of signal-independent noise sources via dark frames. We confirm the inaccuracy of manufacturer DNG noise profiles. We demonstrate how to create and tune a heteroscedastic, physically-based noise model to produce accurate synthesized noise, and use our models to create a large synthesized noisy image dataset, with both RAW and TIFF images. Our noisy images produce excellent results both in LPIPS measurements and in denoising tests.
Our main contributions are as follows:
  • Steps for accurately calibrating different types of cameras, from smartphones to DSLRs, and tuning the resulting heteroscedastic models.
  • Description of iPhone noise suppression for flat-field RAW images.
  • Heteroscedastic noise models that achieve a 54% to 64% improvement in PSNR dB vs. standard DNG noise models.
  • The first public, calibrated dataset providing perfectly aligned noisy images in both RAW and TIFF formats. The dataset covers four cameras, 30 scenes, and over 6000 images in total. The dataset is located on the Harvard Dataverse at https://doi.org/10.7910/DVN/SGHDCP.
  • Source code to calibrate, tune, evaluate, and generate high-quality noisy RAW and TIFF images. The code can be found on GitHub at https://github.com/nikbhatt-cu/SNIC.
3 Related Work
In this section, we will discuss the principal ways in which noise models and datasets are created to train machine learning denoising models. While learning-based noise generation methods exist [2], this paper focuses on physics-based statistical approaches for dataset creation, and we do not discuss denoising models in depth, of which there are many.
3.1 Additive White Gaussian Noise
Additive White Gaussian Noise (AWGN) is a simple noise model. AWGN adds noise to all pixels, regardless of intensity, using a simple Gaussian distribution with zero mean and a chosen variance [2]:
$n\sim\mathcal{N}(0,\sigma^{2})$
A somewhat improved approach scales the standard deviation by the square root of the desired ISO value relative to the base ISO of the camera, reflecting how sensor noise increases with gain:
$n_{i}\sim\mathcal{N}\left(0,\sigma^{2}\right),\quad\text{where }\sigma=\sigma_{\text{base}}\sqrt{\frac{\text{ISO}}{\text{ISO}_{\text{base}}}}$
While easy to implement, AWGN unrealistically treats noise as purely signal-independent, whereas real noise is both signal-dependent and signal-independent. Figure 1 compares AWGN with real noise at a high ISO value.
Real: ISO 100
Real: ISO 12800
AWGN: ISO 12800
Figure 1: Comparison of real camera noise (Sony A7R III) at base and high ISO versus synthetic AWGN.
3.2 Poisson-Gaussian Noise (P-G)
Poisson-Gaussian (P-G) is a physics-based noise model which improves upon AWGN by including both a signal-dependent component (shot noise), and a signal-independent component (read noise). Shot noise occurs because photon arrival is a quantum random process. For a pixel with mean signal $\mu$, the actual observed photon count includes random photon shot noise $N_{p}$. The total photon count follows a Poisson distribution [1]:
$(N_{p}+\mu)\sim\mathcal{P}(\mu),$
where $\mathcal{P}$ is the Poisson distribution with mean and variance $\mu$. Signal-independent read noise is generated by the sensor’s readout, ADC converters and other circuitry. This noise appears even in the absence of light. Read noise is modeled using a Gaussian distribution with zero mean.
A P-G model includes two terms, the signal-dependent Poisson scaling term ($a$) and the signal-independent Gaussian bias term ($b$). A noisy image $I_{n}$ created using P-G from a clean image $I_{c}$ then follows this formula.
$I_{n}=a\cdot I_{c}+b$
Because creating a complete P-G noise model is painstaking work, researchers typically use preexisting Noise Profiles extracted from DNG RAW images.
3.2.1 DNG Noise Profiles
Adobe, Inc. is the creator of the Digital Negative Specification (DNG) which is an open format for encoding digital images [3]. DNG files can store either RAW or non-RAW images. The DNG NoiseProfile tag (51041) uses the Poisson-Gaussian formulation. There are some weaknesses in the NoiseProfile specification. First, the specification simply contains one or more pairs of numbers (the scaling and bias parameters from Poisson-Gaussian). It does not include the ISO for which the noise applies, and there is no way to provide the parameters for multiple ISO values. Second, while some DNG NoiseProfiles include the parameters for each of the Bayer pattern channels (R, G1, G2, and B), it is very common to find NoiseProfiles that contain only three pairs of numbers (R, G, B), or simply a single pair of numbers which are then applied to all four channels equally (iPhone DNGs contain only this single pair). Since the ISO is not specified in the NoiseProfile tag, one must examine DNG files at different ISOs to determine the actual DNG model.
Finally, beyond the lack of per-channel information, DNG noise profiles provided by camera manufacturers tend to be rather inaccurate, compared to calibrated profiles [1]. We have found that profiles generated by Adobe for proprietary RAW files (such as the Sony ARW format) can be more accurate, but still retain the other issues described above.
3.2.2 Calibration and Dark Current
Modeling noise accurately using a Poisson-Gaussian formulation requires careful calibration (described in Methods). In 1994, [4] discusses calibrating a CCD using a gray card, as well as extracting Fixed Pattern Noise by shifting both the light source and gray card. In our calibrations, we confirmed the finding by [1] that fixed pattern noise is not a relevant component in modern CMOS sensors. Thus, our method for calibration deviates somewhat from [4] and includes a tuning portion using real noisy validation images at a range of ISO levels.
The signal-independent part of noise appears in all images from a given sensor, even completely dark ones (e.g., with a lens cap on). Any electronics component can produce noise [1]; as a result, it is extremely difficult, if not impossible, to fully characterize and model this noise physically. From [1] is the complete formulation for a digital image D:
$D=(K_{a}(I+N_{p}+N_{1})+N_{2}+N_{q})K_{d},$
where $K_{a}$ and $K_{d}$ are the analog and digital gains, $N_{1}$ includes noise affected by analog gain, $N_{2}$ includes thermal and column fixed pattern noise that takes place after the analog gain step, and $N_{q}$ is quantization noise when converting analog values to digital ones.
There are a few techniques for capturing signal-independent noise for the purpose of noise synthesis. The simplest (and most common) is to compute it as part of the linear regression required for P-G noise modeling. The second is to capture dark frames (frames with no light hitting the sensor, such as with a lens cap, and/or in a completely dark room). This second approach, described by [1] as dark current, is used in lieu of the Poisson-Gaussian bias term. It more accurately captures sensor independent noise, as dark images include all noise sources by definition. However, as different sensors (and camera electronics) produce different amounts of noise (and at different stages), we have found that the value of capturing dark current varies substantially from sensor to sensor. Specifically, professional-quality cameras can be very accurately modeled using a standard P-G linear regression. Further, some image signal processors (ISPs) in cameras, such as the iPhone 11 Pro, will process RAW images to suppress or alter noise based on the contents of the scene.
3.3 Denoising Datasets
There are a number of denoising datasets, falling into two categories: ones that use AWGN to synthesize noise, and ones that use real noisy images. Due to the quality issues with AWGN, and the lack of published P-G datasets, we only list real noise datasets. Table 1 lists some popular real noisy image datasets [5], [6], [7], [8], [9], [10], [11].
Dataset Year Images Scenes Cameras Format
DND 2017 50 50 4 DSLRs RAW
SID 2018 5094 424 2 DSLRs RAW
SIDD 2018 30000 10 5 Phones RAW
RENOIR 2018 480 120 3 Mixed RAW
ELD 2020 640 10 4 DSLR RAW
PolyU 2021 80 40 5 DSLR JPEG
MIDD 2024 420000 50 20 Phones PNG
Table 1: Real Noise Datasets. Note: Most RAW datasets listed are from 2017-2018, highlighting the need for more updated datasets.
Real noise datasets group low-ISO clean images with noisy images captured at various ISO values [9]. Perfect alignment is challenging; even sub-millimeter camera movement can cause misalignment. Further, strict lighting requirements limit most datasets to indoor scenes. For example, SIDD contains 30,000 images but only 10 distinct scenes [9], while DND provides just 50 image pairs [8]. Most real noise datasets were created in 2017-2018, predating recent sensor advances. Perfectly aligned real noise is a high bar, but when achieved, it provides an unimpeachable ground truth for training. Some datasets, including [10] and [11], only include JPEG or PNG images, so they are not usable for RAW-based denoising.
Aside from AWGN, the alternative to real noise is physically modeled, Poisson-Gaussian synthesis. DNG RAW images include P-G noise profiles provided by the manufacturer or by Adobe as part of its camera calibration process. Because the quality of P-G models varies greatly, we created our own high-quality noise models. While somewhat laborious to create, our models have advantages over real noise. Our noisy images are perfectly aligned by definition, both indoor and outdoor scenes can be captured with a variety of lighting, and there is no limit to the number of images that can be produced. Further, by using dark current frames for some sensors, complex noise sources can be modeled.
Perhaps due to the difficulty of making high-quality noise models, ours is the first dataset that provides both RAW and 16-bit TIFF images with high-quality synthetic noise already inserted.
4 Methods
4.1 RAW Images and Noise
Most cameras (including the ones we used) use a CMOS chip with a color filter array placed over the sensor sites. The color filter array usually follows the Bayer pattern (also called a mosaic), which arranges red, green, and blue filters in a prescribed order. There are twice as many green pixels as either red or blue pixels, due to the human eye’s increased sensitivity to green. There are two principal ways to process images for noise modeling and synthesis. One is to demosaic the Bayer data into RGB data and measure that data (or simply use TIFF, PNG, or JPEG data). The second is to operate on the Bayer data directly. Our approach works with the Bayer data directly because it provides higher quality. Bayer data is linear, rather than gamma-corrected, and we can ignore white balance and the specifics of the demosaic algorithm or the tone curve applied (of which there are many). Further, the RAW-based noise model can produce different values for each of the two green channels which is not possible once they have been processed into RGB.
4.2 DNG Noise Profile Analysis
The Poisson-Gaussian formulation enables straightforward synthesis of noisy images. However, the quality of this synthesis is entirely dependent on the accuracy and completeness of the formulation. One can either build a noise model manually, or rely on published noise profiles. Because building profiles manually can be labor-intensive and error-prone, most researchers rely on profiles stored as DNG NoiseProfile tags. For proprietary RAW files, the simplest approach is to convert RAW files into DNGs and then read the NoiseProfile built by Adobe’s DNG Converter software. This noise profile is computed by Adobe as part of its process for supporting proprietary RAW files. As mentioned previously, the DNG NoiseProfile specification does not include ISO data, so to build an accurate P-G model, one must provide the DNG converter a series of RAW files, shot with the desired range of ISO values. We found that for two Sony cameras (RX100 IV and A7R III), the Adobe profiles are very similar to the ones we built through our calibration and include noise parameters for red, green, and blue channels (though not for both green channels).
For cameras (including smartphone cameras) that produce DNG files natively, the noise profiles are provided by the manufacturer, not Adobe. We found that the DNG noise profile for the iPhone 11 Pro is both incomplete and inaccurate, especially at higher ISO values, which confirms the findings of [1]. First, the DNG NoiseProfile for the iPhone 11 Pro does not include scaling and offset values for each channel, providing instead only a single pair of values. Second, as the ISO rises, the scaling factor should rise, since noise is related to analog gain (ISO). However, for the iPhone 11 Pro, the values level off. This is discussed further in the next section.
Figure 2: Comparison of DNG Profiles. Note: Y-axis scales differ between cameras due to varying sensor characteristics. iPhone noise parameters exhibit non-linearity at high ISO values.
To verify that iPhone DNG noise profile issues are not unique to the iPhone 11 Pro, we extracted DNG noise profiles from images captured with the iPhone 15 Pro (released in 2023). The DNG noise profiles for the iPhone 15 Pro also only include a single pair of values, rather than per-channel noise parameters. Figure 3 compares the scaling term for both iPhone 11 Pro and iPhone 15 Pro cameras. All four iPhone DNG models exhibit similar non-linear behavior at high ISO values, with the iPhone 15 Pro DNG model showing somewhat greater irregularities. This confirms that DNG noise profile inaccuracies occur in multiple iPhone generations, further pointing to the need for calibration and tuning.
Figure 3: DNG NoiseProfile scaling term (a) for iPhone 11 Pro and iPhone 15 Pro across ISO values. Both models show non-monotonic behavior at high ISOs.
4.3 Non-linearity in iPhone Noise Models
We observed a distinct difference between the noise models for the Sony cameras and the iPhone main and telephoto cameras. While the Sonys had a fairly consistent linear relationship (in log-log space) with ISO, the models for the iPhone sensors sharply level off, which is physically impossible. The Sony noise models were created by Adobe, while the iPhone captures in DNG format already, so the models are written into the original files by Apple. This leveling-off would point to one of two explanations: either the calibration was performed incorrectly, or the images used to compute the noise parameters do not show the expected noise. Because we replicated this same leveling-off in our own calibration process for the iPhone cameras, it pointed to the latter explanation. Visually inspecting RAW DNG images confirmed our suspicions: at the same ISO and exposure time, despite being Bayer pattern RAW images, our flat-field iPhone images do not have the same noise characteristics as validation RAW images containing texture and complexity.
Calibration Image
Validation Image
Figure 4: Crops of calibration and validation iPhone 11 Pro images shot at the same ISO and exposure times.
Thus, to produce iPhone images with accurately synthesized noise, we found our calibrated model must be further tuned using carefully captured validation images at a range of ISO values.
We validated our models using 6 scenes captured across the four cameras (3 scenes each captured with the iPhone and Sony RX100 IV, and 3 additional scenes captured with the Sony A7R III). Each validation scene was captured at a low ISO (clean reference) and at 6-7 noisy ISO levels in powers of two, with two noisy captures per ISO level for baseline noise analysis. This yielded 156 full-resolution RAW validation images spanning most or all of the true ISO range of each camera. All validation images were captured using the same tripod and tethering setup as calibration to minimize motion. Unlike the flat-field calibration images, the validation images contained diverse textures, colors, and content, with a mix of indoor and outdoor scenes.
4.4 Calibration Procedure
We calibrated noise models for the iPhone 11 Pro main camera, iPhone 11 Pro telephoto camera, Sony RX100 IV, and Sony A7R III. The iPhone main sensor has 12 megapixels and a fixed 35mm equivalent of 26mm. The telephoto camera is also 12 megapixels with a fixed 35mm equivalent of 52 mm. The Sony RX100 IV captures 20 megapixel images with an attached zoom lens with an approximate 35mm equivalent range of 24mm to 70mm. For the 42 megapixel full-frame Sony A7R III, we attached the Sony FE 24-105mm F4 G OSS lens. This was also the lens used for the capture of all calibration and validation images. That lens was chosen to ensure a similar focal length range with the iPhone and RX100 IV when capturing dataset images. For the Sony RX100 IV and A7R III, we captured images at a median aperture and focal length to minimize lens distortion and shading.
Building a calibrated and tuned noise model is painstaking work, but the resulting synthesized images are considerably more realistic and accurate. We calibrated the cameras outdoors during clear, sunny days to ensure consistent, even lighting. The first step is taking a series of flat-field images. In order to fill the frame, we attached a large poster board to a sheltered, north-facing wall. We also used the ChromLives White Balance card for some tests. All cameras were mounted on a sturdy Manfrotto tripod and leveled in all three dimensions, and positioned close to the poster or card while taking care to avoid casting shadows on the subject. For the Sony cameras, we used Sony’s Imaging Edge tethering software on a Windows PC to control the camera without touching it. For the iPhone, we developed a custom app to control the iPhone camera. The iPhone app provided full control over the ISO and exposure times, while ensuring the DNGs were as unprocessed as Apple allows. In conjunction, we developed a macOS tethering app to control the iPhone app over a wireless network.
4.4.1 Flat-Field Calibration
For all cameras, we selected the lowest true ISO for the camera, and focused manually to infinity to avoid capturing texture in the poster or card we used. The iPhone camera lenses have fixed apertures; for the Sony cameras, we selected an aperture of f/8 to decrease vignetting and a midpoint focal length to decrease distortion. We also disabled Sony’s in-camera long exposure noise reduction, and recorded uncompressed RAW images, rather than Sony’s default lossy compressed format. For all cameras, we took pairs of RAW images at a range of ISO values and exposure times in order to capture the full range of brightness values. Our code examined those images to identify the area near the center of the image that had the least variance in pixel values for calibration. We targeted the center region to further minimize lens distortion and shading.
Figure 5: Flat-field crop area for Sony A7R III showing low variance in pixel intensity.
After collecting the flat-field images and identifying the crop area, for each ISO, we grouped and sorted the images into pairs by exposure time. The Bayer pattern data was cropped and the channel means were computed. Images that were too dark (mean <= 0.02) or too bright (mean >= 0.98) were discarded to avoid clipped data. We tiled each pair of surviving images. For each tile pair, we computed the tile means and computed the variance by subtracting one tile from its partner. To estimate the Poisson-Gaussian noise parameters, we used a two-stage procedure: first, a HuberRegressor identified and removed outliers from the mean-variance relationship; second, a standard least-squares linear regression on the resulting cleaned data yielded the final Poisson-Gaussian noise model parameters (photon shot noise coefficient $a$ and read noise variance $b$).
We built Photon Transfer Curves (PTC) for all cameras and ISO values using the red channel. Figure 6 shows a representative plot at ISO 3200. The red line indicates the fitted P-G model ($\sigma^{2}{}=a\mu+b$), with R2=0.9995. The parameters $a$ and $b$ from this fit are then used in the P-G synthesis model. The increase in spread of measured points as mean signal rises is consistent with photon shot noise, while the presence of outliers at the higher signal levels demonstrates the need for robust regression techniques.
Figure 6: PTC for iPhone 11 Pro main camera
4.4.2 Dark Current Capture
Read noise is extremely difficult to fully characterize in modern sensors [1]. While flat-field calibration can characterize read noise, a more robust technique uses dark frames. Dark frames are captured with a lens cap when possible, and in a dark room or environment. We placed cameras face down inside black fabric pouches and placed those inside sealed boxes, with only a small cut out for the tethering cable. Then per the technique in [1] we captured 10 dark frames at each ISO at a standard exposure time of 1 second. It is essential to disable any in-camera noise reduction to get accurate dark frame images. Those dark frame images are used in the synthesis process as described below.
4.5 Tuning the Noise Model
After observing non-linearity in the initial noise models for the Sony RX100 IV and the iPhone, we confirmed the issue by visually comparing real noisy images with calibration images, especially at high ISO values where it is most noticeable. To normalize the noise model across ISO settings, we fit a power-law relationship between shot noise parameter $a$ and ISO sensitivity. For each color channel, we constrained the fit to pass exactly through measurements at the 2nd and 4th ISO values (e.g., ISO 50 and 200 for the iPhone), which empirically exhibited high measurement reliability (R² > 0.999). The relationship follows:
$a=k\cdot\text{ISO}^{m}$
where k and m are channel-specific constants determined by the two anchor points. We performed this tuning for the iPhone sensors and for the Sony RX100 IV, which also shows some non-linearity at the highest ISO values. The Sony A7R III required no such scaling. Our tuning improvements resulted in superior LPIPS measurements vs. the images that were calibrated but not tuned. Figure 7 compares the DNG noise model with a model that was only calibrated, and also one that was both calibrated and then tuned as described.
Figure 7: DNG, Calibrated, and Tuned Noise Models
4.6 Noise Synthesis
To generate noisy images, we start with low-ISO RAW images from each camera. These “validation” images have texture and mixed subject matter, rather than uniform, flat-fields. When capturing these low ISO images, we also capture pairs of noisy images at a range of ISO values. All validation images are captured using the same tripod and tethering software used for calibration to minimize motion between captures. For each target noisy ISO, we load the calibrated per-channel noise model and, optionally, the list of dark frame images. Not all cameras benefit from dark current synthesis. Specifically we found the Sony RX100 IV and Sony A7R III noise models were extremely accurate and additional dark current worsened the synthesis quality. For the iPhone cameras, introduction of dark current improved the quality.
Given a clean low-ISO RAW image Bayer pattern channel $I_{\text{clean}}$ normalized to $[0,1]$, we synthesize and inject Poisson-Gaussian noise as follows:
$\lambda_{\text{noisy}}\sim\mathcal{P}\left(\frac{I_{\text{clean}}}{a}\right)$
$I_{\text{noisy}}=a\cdot\lambda_{\text{noisy}}+n_{\text{read}},\quad\text{where }n_{\text{read}}\sim\mathcal{N}(0,b)$
where $a$ is the photon-to-digital-number conversion factor, and $b$ is the read noise variance. This process is applied to all four color channels (R, G1, G2, B) of the RAW Bayer pattern. This approach is used for cameras that do not benefit from dark current synthesis (the Sony cameras in our case).
When including dark current, we do not use the read noise from the noise model. Instead, one of the ten dark frames is randomly selected. Following [1], rather than random sampling from the chosen dark current frame, we inject dark current with spatial alignment maintained to increase the accuracy of the synthesis. Once the final noisy Bayer data was generated, we produced multiple outputs: First, we wrote the noisy Bayer data to a binary file along with image metadata in a separate JSON file. Second, we demosaiced the noisy Bayer data and rendered them to 16-bit TIFF files. Then, we demosaiced the clean RAW and validation RAW images and rendered them to 16-bit TIFFs using the same RAW decoding parameters. For the dataset, we also created noisy RAW images. For those, we developed a C++ command line tool, linked to the Adobe DNG SDK, to read the binary and JSON files written earlier to create standard RAW DNGs.The result is a collection of RAW and TIFF images: clean, real noise, and synthesized noise.
(a) Clean Image
(b) Real Noisy Image
(c) Ours (Calibrated + Tuned)
(d) DNG Noise Model
Figure 8: Comparison of noise at ISO 1600 for the iPhone 11 Pro telephoto camera
5 Results
Our images with noise introduced using calibrated and tuned noise models are more realistic than images using DNG noise models. The improvement in realism is greater for the iPhone cameras and lesser for the Sony RX100 IV and Sony A7R III, but superior in all cases. In addition, we evaluated our physically modeled noise with the NAFNet [12] denoising model, which showed superior comparative results when tested with our images.
5.1 Evaluating Synthesis Quality
To assess synthesis quality, we compared synthesized noisy images to real noisy images. To establish a baseline, we compared pairs of real noisy images captured in succession at each ISO. We considered three metrics: LPIPS [13] (lower is better), SSIM [14] (higher is better), and PSNR (higher is better). While PSNR would seem to be the obvious metric, it is only suitable for comparing clean images to denoised images. To confirm that, we computed the PSNR for the iPhone 11 Pro main camera, using a set of real noisy image pairs. The PSNR value ranged from -69 dB to below -80 dB. The same issue affects the SSIM metric. When used to compare two real noisy images, the SSIM metric approached 0.0 (no similarity). In contrast, LPIPS, a learned perceptual similarity metric, was a more effective metric. A low LPIPS score indicates greater similarity between two images, and unlike PSNR or SSIM, this metric consistently identified the pairs of real noisy images as similar, rather than dissimilar. When comparing different synthesis techniques, we computed the LPIPS value at each ISO for a real noisy image pair, and then plotted the absolute value of the difference between that value and the LPIPS value of a pair of synthesized noisy image and a real noisy image. We also plotted absolute LPIPS values for real noisy pairs and pairs of synthesized noisy images and real noisy images.
(e) PSNR
(f) SSIM
(g) LPIPS
Figure 9: iPhone 11 Pro Metrics for Real Noisy Image Pairs
5.2 LPIPS results
Table 2: Quantitative results from NAFNet (PSNR/SSIM) comparing synthesis methods across cameras. Real-world noise serves as the baseline. DNG Model represents Poisson-Gaussian models from DNG noise profiles (manufacturer or Adobe-provided). Smaller gaps indicate more realistic synthesis. Best synthesis methods per metric are emboldened and highlighted in blue.
Camera Metric Real DNG Model Ours Gap (DNG) Gap (Ours) Improvement
iPhone 11 Pro Main (10,368 pairs) PSNR 25.91 31.24 27.80 5.33 1.89 64.5%
SSIM 0.799 0.940 0.882 0.141 0.083 41.1%
iPhone 11 Pro Telephoto (8,640 pairs) PSNR 27.48 31.21 29.21 3.73 1.73 53.6%
SSIM 0.850 0.936 0.899 0.086 0.049 43.0%
Sony RX100 IV (17,820 pairs) PSNR 32.57 32.93 32.86 0.36 0.29 19.4%
SSIM 0.911 0.913 0.911 0.002 0.000 100.0%
Sony A7R III (42,336 pairs) PSNR 36.59 37.95 37.90 1.36 1.31 3.7%
SSIM 0.961 0.966 0.966 0.005 0.005 0.0%
We have provided plots for the four camera sensors, using both absolute and relative measurements. The left plot shows the LPIPS absolute value (lower is better) for each combination of Real Noisy 1, Real Noisy 2, DNG model, and ours. The right plot shows a reduced chart that is relative to the Real Noisy 1 / Real Noisy 2 pair, where closer to zero is best. The calibrated and tuned noise for the two iPhone cameras are clearly superior to the DNG model noise and close to the real noisy images, even at higher ISO values.
(a) Absolute LPIPS
(b) Relative LPIPS
Figure 10: iPhone 11 Pro main camera
Figure 8 shows full-size crops of real noise, DNG model noise, and our tuned and calibrated noise for the iPhone 11 Pro telephoto lens at ISO 1600.
5.3 NAFNet Denoising Results
The Nonlinear Activation Free Network (NAFNet) [12] is a simple denoising model that, despite its simplicity, achieves SOTA results. We used their model pretrained on the real-noise SIDD dataset (NAFNet-SIDD-width64) [6]. We constructed LMDB datasets for each camera with three noise approaches: real noise, the DNG model, and our calibrated and tuned model. Each was paired with the low-ISO ground truth images.
Because the SIDD validation set uses 256×256 pixel patches, we generated 256×256 patches from our full-resolution validation images. Our validation set comprises 79164 patch pairs — over 60× more than SIDD’s 1280 validation patches. The iPhone cameras contributed over 8000 patches each, while higher-resolution Sony cameras produced over 17000 patches each.
This experiment evaluates how well NAFNet generalizes to different noise types (real, DNG model, and ours) using identical scenes, cameras, and ISO values, rather than comparing absolute performance with SIDD. Using PSNR and SSIM metrics, NAFNet denoised our synthesized noisy images significantly more accurately than noisy images created with DNG models. For the iPhone main camera, the PSNR gap between our synthesized noise and real noise was 1.89 dB versus 5.33 dB for DNG model synthesis — a 64.5% improvement. The SSIM gap was 0.083 versus 0.141, a 41.1% improvement. These results demonstrate that our calibrated models produce substantially more realistic noise than DNG models. Complete results appear in Table 2.
6 Dataset
In addition to building accurate heteroscedastic noise models, we have created the first public dataset of RAW images with synthesized noise. Images and models are from four camera sensors: iPhone 11 Pro Main, iPhone 11 Pro Telephoto, Sony RX100 IV, and Sony A7R III. To match the iPhone focal lengths, we captured images for the Sony RX100 IV and A7R III at approximately 24 and 50mm (equivalent). We also captured additional images with the two Sonys at 70mm and 105mm respectively.
Table 3: Synthesized Noisy Images using Calibration dataset (SNIC)
Camera Originals Total Images Total Size (GB)
iPhone 11 Pro Main 132 1848 79.28
iPhone 11 Pro Telephoto 143 2002 85.68
Sony RX100 IV 79 1264 92.47
Sony A7R III 93 1488 229.63
Totals 447 6602 487.06
For each sensor, we captured the same 30 scenes (15 indoor and 15 outdoor). The scenes cover a range of lighting, colors, and textures. All images were captured in RAW format at the sensor’s lowest real ISO. Then, we injected noise for a range of ISO values using our calibrated and tuned noise models into the low-ISO RAW original, using the dark frame database when appropriate. For the dataset, for each camera and scene, we included the low-ISO RAW image and a 16-bit TIFF image rendered from that RAW image. Then, we included both a noisy RAW DNG and a noisy 16-bit TIFF at a range of ISO values. The dataset contains approximately 6600 images. The dataset is located on the Harvard Dataverse at https://doi.org/10.7910/DVN/SGHDCP.
7 Conclusion
In this paper, we describe the advantages of creating calibrated and tuned noise models for synthesizing noisy images, in conjunction with dark current frames. Our approach produces images that are quantitatively superior to both manufacturer and Adobe noise models. We examined our images visually, used the LPIPS metric, and tested our images with the NAFNet denoising model. In each case, our images outperformed the published DNG Poisson-Gaussian models, often dramatically so. Our models are created per ISO and per channel, using carefully captured calibration images, and sound mathematical modeling. The result is the first synthetic noise dataset composed of both RAW and TIFF images, from four camera sensors and with a diverse collection of indoor and outdoor scenes. Researchers can use our publicly available dataset for training, or use our code to create their own.
Appendix A Appendix
Following are additional plots of LPIPS results. The “Absolute LPIPS” figures score pairs of images in absolute terms, while the “Relative LPIPS” figures score each synthesis technique against a zero-baseline of the pairs of real noisy images. In all cases for these LPIPS figures, a lower score is better. Following the figures are comparisons of real and synthesized noisy images from each camera.
(a) iPhone 11 Pro Main
(b) iPhone 11 Pro Telephoto
Figure 11: iPhone LPIPS measurements
(a) Sony RX100 IV
(b) Sony A7R III
Figure 12: Sony LPIPS measurements
(a) Clean Image
(b) Real Noisy Image
(c) Ours (Tuned and Calibrated)
(d) DNG
Figure 13: Comparison of noise at ISO 800 for the iPhone 11 Pro main camera
(a) Clean Image
(b) Real Noisy Image
(c) Ours (Tuned and Calibrated)
(d) DNG
Figure 14: Comparison of noise at ISO 12800 for the Sony RX100 IV
(a) Clean Image
(b) Real Noisy Image
(c) Ours (Tuned and Calibrated)
(d) DNG
Figure 15: Comparison of noise at ISO 6400 for the Sony A7R III

### v2

@@
 SNIC: Synthesized Noisy Images using Calibration
-SNIC: Synthesized Noisy Images using Calibration
-Nik Bhatt
-Columbia University
-nb3097@columbia.edu
-1 Abstract
+Abstract
 Advanced denoising algorithms require large, high-quality datasets.
-Physics-based, statistical noise models can create such datasets by realistically simulating noise in digital images.
+Physically-based statistical noise models can create such datasets by realistically simulating noise in digital images.
 However, there is little information on the correct way to calibrate and tune these heteroscedastic models, and a lack of published datasets using them.
 In this paper, we explore the process of building high-quality heteroscedastic noise models.
-Our methods produce realistic synthesized noisy images, in both RAW and TIFF formats.
-Our synthesized noisy images achieve comparable LPIPS results to real noisy images, and greatly outperform those created with manufacturer-provided DNG noise models both in LPIPS and when tested with a state-of-the-art (SOTA) denoising model.
+Our methods produce realistic synthesized noisy images in both RAW and TIFF formats.
+Our synthesized noisy images achieve comparable LPIPS results to real noisy images; when tested with a state-of-the-art denoising model, our images reduce the PSNR gap versus real noise by 54–64% compared to those synthesized using manufacturer-provided DNG noise models.
 Using our approach, we created the Synthesized Noisy Images using Calibration dataset (SNIC) containing over 6000 noisy images, comprising 30 scenes from four sensors, including two smartphone sensors, a point-and-shoot, and a DSLR.
 SNIC is the first synthesized noisy image dataset provided in both RAW and TIFF format.
+Keywords: Denoising, Noise Modeling, Poisson-Gaussian, Image Dataset, RAW, Smartphone, Calibration
 1 Introduction
@@
 Noise arises from the quantum effects of photon arrival, and from various electronic components in the sensor and camera circuitry, including sensor readout, amplifiers, and analog-to-digital conversion.
-Images captured in low-light environments, and especially with the smaller sensors found on smartphones, typically include considerable amounts of noise, negatively affecting the appearance and aesthetics of the images.
+Images captured in low-light environments, and especially with the smaller sensors found on smartphones, typically include considerable amounts of noise.
+This negatively affects the appearance and aesthetics of the images.
 Thus, noise reduction, or denoising, is an important and difficult task that has been the focus of active research for years.
@@
 Denoising is best performed on RAW images (rather than camera-generated JPEGs) because RAW images are scene-referred, undergo less processing by the image signal processor (ISP), contain linear data, and do not include gamma correction.
-Currently most research in denoising focuses on machine learning models.
+Currently, most research in denoising focuses on machine learning models.
 Thus, researchers developing new denoising algorithms rely on large image datasets to train and test their models.
-There are three main approaches to acquiring an adequate dataset.
+There are three main approaches to acquire an adequate dataset.
   • First, one can add noise to images from another dataset such as Flickr2K.
@@
 Consequently, AWGN-based synthetic images lack realism, failing to capture the true noise characteristics of digital sensors.
-  • Second, one can use an existing real noisy dataset that consists of low ISO and high ISO images that have been extremely carefully aligned.
+  • Second, one can use an existing real noisy dataset that consists of low- and high-ISO images that have been extremely carefully aligned.
 Making a real noise dataset is exceptionally difficult, so there are only a few available.
-Most were created in 2017 or 2018, and thus use older cameras such as the iPhone 7, released over nine years ago.
+Most were created in 2017 or 2018, and thus use older cameras such as the iPhone 7, released more than nine years ago.
 Real noise datasets are also often limited to specially lit indoor scenes.
@@
 A heteroscedastic model can be easily retrieved from DNG images, though we found those models are often inaccurate and incomplete.
-We have chosen a variant of the third approach: create a custom physics-based heteroscedastic noise model by calibrating cameras and then tuning the model with validation images.
-This requires more effort than simply using DNG noise profiles, but our models provide perfectly aligned, high-quality noisy RAW images.
+We have chosen a variant of the third approach: create a custom physically-based heteroscedastic noise model by calibrating cameras and then refining the noise model with special tuning images.
+We use those models to introduce noise into clean reference RAW images.
+This approach requires more effort than simply using DNG noise profiles, but our models provide perfectly aligned, high-quality noisy RAW images.
 Further, this approach offers full control over the choice of cameras and scenes used.
-The choice and quality of the dataset directly affects the performance of a denoising algorithm on real images.
-For physically modeled datasets, the key is minimizing the discrepancy between the synthetic noise and actual noise [1].
+The choice and quality of the training dataset directly affects the performance of a denoising model on real images.
+For physically modeled datasets, the key is minimizing the discrepancy between the synthetic noise and actual noise (Zhang et al., 2021).
 In this paper, we focus on physics-based noise modeling, using careful calibration to determine photon shot noise, combined with complete capture of signal-independent noise sources via dark frames.
@@
   • Description of iPhone noise suppression for flat-field RAW images.
-  • Heteroscedastic noise models that achieve a 54% to 64% improvement in PSNR dB vs. standard DNG noise models.
+  • Heteroscedastic noise models that achieve a 54% to 64% reduction in PSNR gap vs. standard DNG noise models.
   • The first public, calibrated dataset providing perfectly aligned noisy images in both RAW and TIFF formats.
@@
 2 Related Work
-In this section, we will discuss the principal ways in which noise models and datasets are created to train machine learning denoising models.
-While learning-based noise generation methods exist [2], this paper focuses on physics-based statistical approaches for dataset creation, and we do not discuss denoising models in depth, of which there are many.
+This section discusses the principal ways in which noise models and datasets are created to train machine learning denoising models.
+While learning-based noise generation methods exist (Abdelrahman et al., 2019), this paper focuses on physics-based statistical approaches for dataset creation, and we do not discuss denoising models in depth, of which there are many.
 2.1 Additive White Gaussian Noise
-Additive White Gaussian Noise (AWGN) is a simple noise model. AWGN adds noise to all pixels, regardless of intensity, using a simple Gaussian distribution with zero mean and a chosen variance [2]:
+Additive White Gaussian Noise (AWGN) is a simple noise model.
+AWGN adds noise to all pixels, regardless of intensity, using a simple Gaussian distribution with zero mean and a chosen variance (Abdelrahman et al., 2019):
 $n\sim\mathcal{N}(0,\sigma^{2})$
@@
 For a pixel with mean signal $\mu$, the actual observed photon count includes random photon shot noise $N_{p}$.
-The total photon count follows a Poisson distribution [1]:
+The total photon count follows a Poisson distribution (Zhang et al., 2021):
 $(N_{p}+\mu)\sim\mathcal{P}(\mu),$
@@
 Adobe, Inc.
-is the creator of the Digital Negative Specification (DNG) which is an open format for encoding digital images [3].
+is the creator of the Digital Negative Specification (DNG) which is an open format for encoding digital images (Adobe Systems Incorporated, 2023).
 DNG files can store either RAW or non-RAW images.
@@
 Since the ISO is not specified in the NoiseProfile tag, one must examine DNG files at different ISOs to determine the actual DNG model.
-Finally, beyond the lack of per-channel information, DNG noise profiles provided by camera manufacturers tend to be rather inaccurate, compared to calibrated profiles [1].
+Finally, beyond the lack of per-channel information, DNG noise profiles provided by camera manufacturers tend to be rather inaccurate, compared to calibrated profiles (Zhang et al., 2021).
 We have found that profiles generated by Adobe for proprietary RAW files (such as the Sony ARW format) can be more accurate, but still retain the other issues described above.
@@
 Modeling noise accurately using a Poisson-Gaussian formulation requires careful calibration (described in Methods).
-In 1994, [4] discusses calibrating a CCD using a gray card, as well as extracting Fixed Pattern Noise by shifting both the light source and gray card.
-In our calibrations, we confirmed the finding by [1] that fixed pattern noise is not a relevant component in modern CMOS sensors.
-Thus, our method for calibration deviates somewhat from [4] and includes a tuning portion using real noisy validation images at a range of ISO levels.
+Healey and Kondepudy (1994) discuss calibrating a CCD using a gray card, as well as extracting Fixed Pattern Noise by shifting both the light source and gray card.
+In our calibrations, we confirmed the finding by Zhang et al.
+(2021) that fixed pattern noise is not a relevant component in modern CMOS sensors.
+Thus, our method for calibration deviates somewhat from Healey and Kondepudy (1994) and includes a tuning portion using real noisy images at a range of ISO levels.
 The signal-independent part of noise appears in all images from a given sensor, even completely dark ones (e.g., with a lens cap on).
-Any electronics component can produce noise [1]; as a result, it is extremely difficult, if not impossible, to fully characterize and model this noise physically.
-From [1] is the complete formulation for a digital image D:
+Any electronics component can produce noise (Zhang et al., 2021); as a result, it is extremely difficult, if not impossible, to fully characterize and model this noise physically.
+Zhang et al.
+(2021) provide the complete formulation for a digital image D:
 $D=(K_{a}(I+N_{p}+N_{1})+N_{2}+N_{q})K_{d},$
-where $K_{a}$ and $K_{d}$ are the analog and digital gains, $N_{1}$ includes noise affected by analog gain, $N_{2}$ includes thermal and column fixed pattern noise that takes place after the analog gain step, and $N_{q}$ is quantization noise when converting analog values to digital ones.
+where $K_{a}$ and $K_{d}$ represent the analog and digital gains, $N_{1}$ includes noise affected by analog gain, $N_{2}$ includes thermal and column fixed pattern noise that takes place after the analog gain step, and $N_{q}$ is quantization noise when converting analog values to digital ones.
 There are a few techniques for capturing signal-independent noise for the purpose of noise synthesis.
@@
 The second is to capture dark frames (frames with no light hitting the sensor, such as with a lens cap, and/or in a completely dark room).
-This second approach, described by [1] as dark current, is used in lieu of the Poisson-Gaussian bias term.
-It more accurately captures sensor independent noise, as dark images include all noise sources by definition.
+This second approach, described by Zhang et al.
+(2021) as dark current, is used in lieu of the Poisson-Gaussian bias term.
+It more accurately captures sensor-independent noise, as dark images include all noise sources by definition.
 However, as different sensors (and camera electronics) produce different amounts of noise (and at different stages), we have found that the value of capturing dark current varies substantially from sensor to sensor.
@@
 2.3 Denoising Datasets
-There are a number of denoising datasets, falling into two categories: ones that use AWGN to synthesize noise, and ones that use real noisy images.
+There are a number of published denoising datasets, falling into two categories: ones that use AWGN to synthesize noise, and ones that use real noisy images.
 Due to the quality issues with AWGN, and the lack of published P-G datasets, we only list real noise datasets.
-Table 1 lists some popular real noisy image datasets [5], [6], [7], [8], [9], [10], [11].
+Table 1 lists some popular real noisy image datasets (Chen et al., 2018; Abdelhamed et al., 2018; Anaya and Ortiz, 2018; Plötz and Roth, 2017; Wei et al., 2020; Flepp et al., 2024; Xu et al., 2018).
 Dataset Year Images Scenes Cameras Format
@@
 MIDD 2024 420000 50 20 Phones PNG
-Table 1: Real Noise Datasets. Note: Most RAW datasets listed are from 2017-2018, highlighting the need for more updated datasets.
-Real noise datasets group low-ISO clean images with noisy images captured at various ISO values [9].
+Table 1: Real Noise Datasets. Note: Most RAW datasets listed are from 2017-2018, highlighting a need for updated datasets.
+Real noise datasets group low-ISO clean images with noisy images captured at various ISO values (Wei et al., 2020).
 Perfect alignment is challenging; even sub-millimeter camera movement can cause misalignment.
 Further, strict lighting requirements limit most datasets to indoor scenes.
-For example, SIDD contains 30,000 images but only 10 distinct scenes [9], while DND provides just 50 image pairs [8].
+For example, SIDD contains 30,000 images but only 10 distinct scenes (Wei et al., 2020), while DND provides just 50 image pairs (Plötz and Roth, 2017).
 Most real noise datasets were created in 2017-2018, predating recent sensor advances.
 Perfectly aligned real noise is a high bar, but when achieved, it provides an unimpeachable ground truth for training.
-Some datasets, including [10] and [11], only include JPEG or PNG images, so they are not usable for RAW-based denoising.
+Some datasets, including MIDD (Flepp et al., 2024) and PolyU (Xu et al., 2018), only include JPEG or PNG images, so they are not usable for RAW-based denoising.
 Aside from AWGN, the alternative to real noise is physically modeled, Poisson-Gaussian synthesis.
@@
 Our noisy images are perfectly aligned by definition, both indoor and outdoor scenes can be captured with a variety of lighting, and there is no limit to the number of images that can be produced.
-Further, by using dark current frames for some sensors, complex noise sources can be modeled.
+Further, by using dark current frames when appropriate, complex noise sources can be modeled.
 Perhaps due to the difficulty of making high-quality noise models, ours is the first dataset that provides both RAW and 16-bit TIFF images with high-quality synthetic noise already inserted.
@@
 There are two principal ways to process images for noise modeling and synthesis.
-One is to demosaic the Bayer data into RGB data and measure that data (or simply use TIFF, PNG, or JPEG data).
+One is to demosaic the Bayer data into RGB data and measure that data (or simply use pre-rendered TIFF, PNG, or JPEG data).
 The second is to operate on the Bayer data directly.
@@
 Because building profiles manually can be labor-intensive and error-prone, most researchers rely on profiles stored as DNG NoiseProfile tags.
-For proprietary RAW files, the simplest approach is to convert RAW files into DNGs and then read the NoiseProfile built by Adobe’s DNG Converter software.
+For proprietary RAW files, the simplest approach is to convert RAW files into DNGs and then read the NoiseProfile tag written by Adobe’s DNG Converter software.
 This noise profile is computed by Adobe as part of its process for supporting proprietary RAW files.
@@
 For cameras (including smartphone cameras) that produce DNG files natively, the noise profiles are provided by the manufacturer, not Adobe.
-We found that the DNG noise profile for the iPhone 11 Pro is both incomplete and inaccurate, especially at higher ISO values, which confirms the findings of [1].
+We found that the DNG noise profile for the iPhone 11 Pro is both incomplete and inaccurate, especially at higher ISO values, which confirms the findings of Zhang et al.
+(2021).
 First, the DNG NoiseProfile for the iPhone 11 Pro does not include scaling and offset values for each channel, providing instead only a single pair of values.
@@
 This confirms that DNG noise profile inaccuracies occur in multiple iPhone generations, further pointing to the need for calibration and tuning.
-Figure 3: DNG NoiseProfile scaling term (a) for iPhone 11 Pro and iPhone 15 Pro across ISO values. Both models show non-monotonic behavior at high ISOs.
+Figure 3: DNG NoiseProfile scaling term (a) for iPhone 11 Pro and iPhone 15 Pro across ISO values. Both models show unexpected non-monotonic behavior at high ISOs.
 3.3 Non-linearity in iPhone Noise Models
@@
 Because we replicated this same leveling-off in our own calibration process for the iPhone cameras, it pointed to the latter explanation.
-Visually inspecting RAW DNG images confirmed our suspicions: at the same ISO and exposure time, despite being Bayer pattern RAW images, our flat-field iPhone images do not have the same noise characteristics as validation RAW images containing texture and complexity.
+Visually inspecting RAW DNG images confirmed our suspicions: at the same ISO and exposure time, despite being Bayer pattern RAW images, our flat-field iPhone images do not have the same noise characteristics as real noisy RAW images containing texture and complexity.
 Calibration Image
-Validation Image
-Figure 4: Crops of calibration and validation iPhone 11 Pro images shot at the same ISO and exposure times.
-Thus, to produce iPhone images with accurately synthesized noise, we found our calibrated model must be further tuned using carefully captured validation images at a range of ISO values.
-We validated our models using 6 scenes captured across the four cameras (3 scenes each captured with the iPhone and Sony RX100 IV, and 3 additional scenes captured with the Sony A7R III).
-Each validation scene was captured at a low ISO (clean reference) and at 6-7 noisy ISO levels in powers of two, with two noisy captures per ISO level for baseline noise analysis.
-This yielded 156 full-resolution RAW validation images spanning most or all of the true ISO range of each camera.
-All validation images were captured using the same tripod and tethering setup as calibration to minimize motion.
-Unlike the flat-field calibration images, the validation images contained diverse textures, colors, and content, with a mix of indoor and outdoor scenes.
+Tuning Image
+Figure 4: Crops of calibration and tuning iPhone 11 Pro images shot at the same ISO and exposure times.
+Thus, to produce iPhone images with accurately synthesized noise, we found our calibrated model must be further adjusted using carefully captured ”tuning images” at a range of ISO values.
+We tuned and evaluated our models using 6 scenes captured across the four cameras (3 scenes each captured with the iPhone and Sony RX100 IV, and 3 additional scenes captured with the Sony A7R III).
+Each scene was captured at a low ISO (clean reference) and at 6-7 noisy ISO levels in powers of two, with two noisy captures per ISO level for baseline noise analysis.
+This yielded 156 full-resolution RAW tuning images spanning most or all of the true ISO range of each camera.
+All tuning images were captured using the same tripod and tethering setup as calibration to minimize motion (described in Calibration Procedure).
+Unlike the flat-field calibration images, the tuning images contained diverse textures, colors, and content, with a mix of indoor and outdoor scenes.
 3.4 Calibration Procedure
@@
 For the 42 megapixel full-frame Sony A7R III, we attached the Sony FE 24-105mm F4 G OSS lens.
-This was also the lens used for the capture of all calibration and validation images.
+This was also the lens used for the capture of all calibration, tuning, and dataset images.
 That lens was chosen to ensure a similar focal length range with the iPhone and RX100 IV when capturing dataset images.
@@
 We targeted the center region to further minimize lens distortion and shading.
-Figure 5: Flat-field crop area for Sony A7R III showing low variance in pixel intensity.
+Figure 5: Flat-field crop (512x512 pixels) for Sony A7R III showing low variance in pixel intensity.
 After collecting the flat-field images and identifying the crop area, for each ISO, we grouped and sorted the images into pairs by exposure time.
 The Bayer pattern data was cropped and the channel means were computed.
-Images that were too dark (mean <= 0.02) or too bright (mean >= 0.98) were discarded to avoid clipped data.
+Images that were too dark ($\text{mean}\leq 0.02$) or too bright ($\text{mean}\geq 0.98$) were discarded to avoid clipped data.
 We tiled each pair of surviving images.
@@
 3.4.2 Dark Current Capture
-Read noise is extremely difficult to fully characterize in modern sensors [1].
+Read noise is extremely difficult to fully characterize in modern sensors (Zhang et al., 2021).
 While flat-field calibration can characterize read noise, a more robust technique uses dark frames.
 Dark frames are captured with a lens cap when possible, and in a dark room or environment.
-We placed cameras face down inside black fabric pouches and placed those inside sealed boxes, with only a small cut out for the tethering cable.
-Then per the technique in [1] we captured 10 dark frames at each ISO at a standard exposure time of 1 second.
-It is essential to disable any in-camera noise reduction to get accurate dark frame images.
+We placed cameras face down inside black fabric pouches and placed those inside sealed boxes, with a small hole cut for the tethering cable.
+Then per the technique in Zhang et al.
+(2021), we captured 10 dark frames at each ISO at a standard exposure time of 1 second.
+We took care to disable any in-camera noise reduction to get accurate dark frame images.
 Those dark frame images are used in the synthesis process as described below.
@@
 To normalize the noise model across ISO settings, we fit a power-law relationship between shot noise parameter $a$ and ISO sensitivity.
-For each color channel, we constrained the fit to pass exactly through measurements at the 2nd and 4th ISO values (e.g., ISO 50 and 200 for the iPhone), which empirically exhibited high measurement reliability (R² > 0.999).
+For each color channel, we constrained the fit to pass exactly through measurements at the 2nd and 4th ISO values (e.g., ISO 50 and 200 for the iPhone), which empirically exhibited high measurement reliability ($R^{2}>0.999$).
 The relationship follows:
@@
 3.6 Noise Synthesis
-To generate noisy images, we start with low-ISO RAW images from each camera.
-These “validation” images have texture and mixed subject matter, rather than uniform, flat-fields.
-When capturing these low ISO images, we also capture pairs of noisy images at a range of ISO values.
-All validation images are captured using the same tripod and tethering software used for calibration to minimize motion between captures.
-For each target noisy ISO, we load the calibrated per-channel noise model and, optionally, the list of dark frame images.
+We employed noise synthesis at two different stages.
+The first was after calibration and calculation of the initial noise models.
+We captured a series of ”tuning images,” which are sets of clean and real noisy images (described above in Non-linearity in iPhone Noise Models).
+These tuning images were also used to evaluate synthesis quality, as described in the Results section.
+Once the noise models were tuned and evaluated, we created the SNIC dataset by adding synthesized noise to a separate set of clean low-ISO dataset images.
+For each target noisy ISO, we load the calibrated per-channel (and per-ISO) noise model and, optionally, the list of dark frame images.
 Not all cameras benefit from dark current synthesis.
-Specifically we found the Sony RX100 IV and Sony A7R III noise models were extremely accurate and additional dark current worsened the synthesis quality.
+Specifically, we found the Sony RX100 IV and Sony A7R III noise models were extremely accurate and additional dark current worsened the synthesis quality.
 For the iPhone cameras, introduction of dark current improved the quality.
@@
 Instead, one of the ten dark frames is randomly selected.
-Following [1], rather than random sampling from the chosen dark current frame, we inject dark current with spatial alignment maintained to increase the accuracy of the synthesis.
+Following Zhang et al.
+(2021), rather than random sampling from the chosen dark current frame, we inject dark current with spatial alignment maintained to increase the accuracy of the synthesis.
 Once the final noisy Bayer data was generated, we produced multiple outputs: First, we wrote the noisy Bayer data to a binary file along with image metadata in a separate JSON file.
 Second, we demosaiced the noisy Bayer data and rendered them to 16-bit TIFF files.
-Then, we demosaiced the clean RAW and validation RAW images and rendered them to 16-bit TIFFs using the same RAW decoding parameters.
-For the dataset, we also created noisy RAW images.
-For those, we developed a C++ command line tool, linked to the Adobe DNG SDK, to read the binary and JSON files written earlier to create standard RAW DNGs.The result is a collection of RAW and TIFF images: clean, real noise, and synthesized noise.
+Then, we demosaiced both clean and noisy RAW images and rendered each image to a 16-bit TIFF using the same RAW decoding parameters.
+For the dataset, we also created noisy RAW images from the clean RAWs.
+For those, we developed a C++ command line tool, linked to the Adobe DNG SDK, to read the binary and JSON files written earlier to create standard RAW DNGs.
+The result is a collection of clean and synthesized noisy images in both RAW and TIFF formats.
 (a) Clean Image
 (b) Real Noisy Image
-(c) Ours (Calibrated + Tuned)
+(c) Ours (Tuned)
 (d) DNG Noise Model
@@
 The improvement in realism is greater for the iPhone cameras and lesser for the Sony RX100 IV and Sony A7R III, but superior in all cases.
-In addition, we evaluated our physically modeled noise with the NAFNet [12] denoising model, which showed superior comparative results when tested with our images.
+In addition, we evaluated our physically modeled noise with the NAFNet denoising model (Chen et al., 2022), which showed superior comparative results when tested with our images.
+Figure 8 shows full-size crops of real noise, DNG model noise, and our tuned and calibrated noise for the iPhone 11 Pro telephoto lens at ISO 1600.
 4.1 Evaluating Synthesis Quality
 To assess synthesis quality, we compared synthesized noisy images to real noisy images.
+We used the tuning images from the first noise synthesis stage, and added synthesized noise to the low-ISO tuning images using the final tuned noise models.
 To establish a baseline, we compared pairs of real noisy images captured in succession at each ISO.
-We considered three metrics: LPIPS [13] (lower is better), SSIM [14] (higher is better), and PSNR (higher is better).
+We considered three metrics: LPIPS (Zhang et al., 2018) (lower is better), SSIM (Wang et al., 2004) (higher is better), and PSNR (higher is better).
 While PSNR would seem to be the obvious metric, it is only suitable for comparing clean images to denoised images.
 To confirm that, we computed the PSNR for the iPhone 11 Pro main camera, using a set of real noisy image pairs.
-The PSNR value ranged from -69 dB to below -80 dB.
+The PSNR value for real noisy image pairs ranged from -69 dB to below -80 dB.
 The same issue affects the SSIM metric.
 When used to compare two real noisy images, the SSIM metric approached 0.0 (no similarity).
-In contrast, LPIPS, a learned perceptual similarity metric, was a more effective metric.
+In contrast, LPIPS, a learned perceptual similarity metric, was more effective.
 A low LPIPS score indicates greater similarity between two images, and unlike PSNR or SSIM, this metric consistently identified the pairs of real noisy images as similar, rather than dissimilar.
@@
 We also plotted absolute LPIPS values for real noisy pairs and pairs of synthesized noisy images and real noisy images.
+Figure 9 illustrates this for the iPhone 11 Pro camera.
 (e) PSNR
@@
 4.2 LPIPS results
+We have provided plots for the four camera sensors, using both absolute and relative measurements.
+The left plot shows the LPIPS absolute value (lower is better) for each combination of Real Noisy 1, Real Noisy 2, DNG model, and ours.
+The right plot shows a reduced chart that is relative to the Real Noisy 1 / Real Noisy 2 pair, where closer to zero is best.
+The calibrated and tuned noise for the two iPhone cameras are clearly superior to the DNG model noise and close to the real noisy images, even at higher ISO values.
+Figure 10 shows LPIPS results for the iPhone 11 Pro main camera.
+(a) Absolute LPIPS
+(b) Relative LPIPS
+Figure 10: iPhone 11 Pro main camera
+4.3 NAFNet Denoising Results
+The Nonlinear Activation Free Network (NAFNet) (Chen et al., 2022) is a denoising model that, despite its simplicity, achieves state-of-the-art results.
+We used their model pretrained on the real-noise SIDD dataset (NAFNet-SIDD-width64) (Abdelhamed et al., 2018).
+We constructed LMDB datasets for each camera with three noise approaches: real noise, the DNG model, and our calibrated and tuned model.
+Each was paired with the low-ISO ground truth images.
+[…]
+The purpose of this experiment was to evaluate how well NAFNet generalizes to different noise types (real, DNG model, and ours) using identical scenes, cameras, and ISO values, rather than comparing absolute performance with SIDD.
+Using PSNR and SSIM metrics, NAFNet denoised our synthesized noisy images significantly more accurately than noisy images created with DNG models.
+For the iPhone main camera, the PSNR gap between our synthesized noise and real noise was 1.89 dB versus 5.33 dB for DNG model synthesis — a 64.5% improvement.
+The SSIM gap was 0.083 versus 0.141, a 41.1% improvement.
+These results demonstrate that our calibrated models produce substantially more realistic noise than DNG models.
 Table 2: Quantitative results from NAFNet (PSNR/SSIM) comparing synthesis methods across cameras.
@@
 SSIM 0.961 0.966 0.966 0.005 0.005 0.0%
-We have provided plots for the four camera sensors, using both absolute and relative measurements.
-The left plot shows the LPIPS absolute value (lower is better) for each combination of Real Noisy 1, Real Noisy 2, DNG model, and ours.
-The right plot shows a reduced chart that is relative to the Real Noisy 1 / Real Noisy 2 pair, where closer to zero is best.
-The calibrated and tuned noise for the two iPhone cameras are clearly superior to the DNG model noise and close to the real noisy images, even at higher ISO values.
-(a) Absolute LPIPS
-(b) Relative LPIPS
-Figure 10: iPhone 11 Pro main camera
-Figure 8 shows full-size crops of real noise, DNG model noise, and our tuned and calibrated noise for the iPhone 11 Pro telephoto lens at ISO 1600.
-5.3 NAFNet Denoising Results
-The Nonlinear Activation Free Network (NAFNet) [12] is a simple denoising model that, despite its simplicity, achieves SOTA results.
-We used their model pretrained on the real-noise SIDD dataset (NAFNet-SIDD-width64) [6].
-We constructed LMDB datasets for each camera with three noise approaches: real noise, the DNG model, and our calibrated and tuned model.
-Each was paired with the low-ISO ground truth images.
-Because the SIDD validation set uses 256×256 pixel patches, we generated 256×256 patches from our full-resolution validation images.
-Our validation set comprises 79164 patch pairs — over 60× more than SIDD’s 1280 validation patches.
-The iPhone cameras contributed over 8000 patches each, while higher-resolution Sony cameras produced over 17000 patches each.
-This experiment evaluates how well NAFNet generalizes to different noise types (real, DNG model, and ours) using identical scenes, cameras, and ISO values, rather than comparing absolute performance with SIDD.
-Using PSNR and SSIM metrics, NAFNet denoised our synthesized noisy images significantly more accurately than noisy images created with DNG models.
-For the iPhone main camera, the PSNR gap between our synthesized noise and real noise was 1.89 dB versus 5.33 dB for DNG model synthesis — a 64.5% improvement.
-The SSIM gap was 0.083 versus 0.141, a 41.1% improvement.
-These results demonstrate that our calibrated models produce substantially more realistic noise than DNG models.
-Complete results appear in Table 2.
 5 Dataset
@@
 We also captured additional images with the two Sonys at 70mm and 105mm respectively.
+These cameras were selected to span a range of camera types (smartphone, point-and-shoot, and DSLR), a range of sensor sizes and capabilities, and for their tethering capabilities.
+In addition, the Sony cameras and lenses were chosen because their focal lengths are similar to those of the fixed lens iPhone cameras.
+Tethering eliminates the need to touch the camera when configuring the camera or capturing images.
+This is important because touching the camera invariably moves it a small amount, affecting alignment and quality, especially at longer exposure times.
+The models provided in the SNIC dataset are specific to these cameras.
+However, the calibration code and techniques are generic and can be used by researchers to create noise models for other cameras.
 Table 3: Synthesized Noisy Images using Calibration dataset (SNIC)
@@
 For each sensor, we captured the same 30 scenes (15 indoor and 15 outdoor).
-The scenes cover a range of lighting, colors, and textures.
+The scenes and images are distinct from the tuning images described in the Noise Synthesis section, and cover a range of locations, lighting, colors, and textures.
+Further, there are no real noisy images captured for the dataset.
 All images were captured in RAW format at the sensor’s lowest real ISO.
@@
 The dataset is located on the Harvard Dataverse at https://doi.org/10.7910/DVN/SGHDCP.
+A datasheet for the dataset is provided as supplementary material.
 6 Conclusion
@@
 Researchers can use our publicly available dataset for training, or use our code to create their own.
-Appendix A Appendix
+Broader Impact Statement
+The SNIC dataset can be used both to train and to test noise reduction models.
+Such models can be used in a wide range of applications, such as image processing and computer vision.
+The calibration methods described and code provided would allow others to create datasets for other cameras, or to validate existing noise models.
+A high quality dataset in RAW and TIFF formats allows researchers to work at different spots in either linear or gamma-corrected space, which offers the potential for more sophisticated noise reduction models to be created and validated.
+The code provides calibration, tuning, and injection scripts to create new high quality datasets.
+Improved noise reduction models would result in higher quality images and videos, especially in low light environments.
+At the same time, noise reduction can be used in computer vision to clean up still images and video frames from video surveillance.
+While the dataset and code do not themselves provide the ability to enhance surveillance, improved denoising could enhance surveillance capabilities through better quality low-light video frames.
+Acknowledgments and Disclosure of Funding
+No external financial support or grants were received from any public, commercial, or not-for-profit entities for the research, authorship, or publication of this article.
+The author declares that he has no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Appendix A Additional Results
 Following are additional plots of LPIPS results.
-The “Absolute LPIPS” figures score pairs of images in absolute terms, while the “Relative LPIPS” figures score each synthesis technique against a zero-baseline of the pairs of real noisy images.
+The “Absolute LPIPS” figures score pairs of tuning images in absolute terms, while the “Relative LPIPS” figures score each synthesis technique against a zero-baseline of the pairs of real noisy images.
 In all cases for these LPIPS figures, a lower score is better.
@@
 (b) iPhone 11 Pro Telephoto
-Figure 11: iPhone LPIPS measurements
-(a) Sony RX100 IV
-(b) Sony A7R III
-Figure 12: Sony LPIPS measurements
+(c) Sony RX100 IV
+(d) Sony A7R III
+Figure 11: LPIPS measurements for all cameras
 (a) Clean Image
 (b) Real Noisy Image
-(c) Ours (Tuned and Calibrated)
+(c) Ours (Tuned)
 (d) DNG
@@
 (b) Real Noisy Image
-(c) Ours (Tuned and Calibrated)
+(c) Ours (Tuned)
 (d) DNG
@@
 (b) Real Noisy Image
-(c) Ours (Tuned and Calibrated)
+(c) Ours (Tuned)
 (d) DNG